OpenStudy (anonymous):

Please help establish the identity. 1-[(sin^2v)/(1+cosv)]= cosv

OpenStudy (solomonzelman):

\[1- \frac{\sin^2v}{1+cosv}=\cos{v}\] \[\frac{1+cosv}{1+cosv}- \frac{\sin^2v}{1+cosv}=\cos{v}\] \[\frac{1+cosv-\sin^2v}{1+cosv}=\cos{v}\] \[\frac{(1-\sin^2v)+cosv}{1+cosv}=\cos{v}\] \[\frac{\cos^2v+cosv}{1+cosv}=\cos{v}\] \[\frac{cosv(cosv+1)}{(1+cosv)}=\cos{v}\]got it from here ?

OpenStudy (anonymous):

amazing!

OpenStudy (solomonzelman):

Tnx, my best class 98 in trig :))))

OpenStudy (solomonzelman):

What about the previous one, do you need help with it ?